OpenStudy (raden):

integral of dx/(2 + 2 cos(x) - 2sqrt(3)sin(x))

OpenStudy (mathmale):

You have 2 terms in this integrand, separated by a negative sign. The two terms have no relationship to each other that I can see. So your best bet may be to attempt to integrate each of the 2 terms separately. What is Int(dx/(2 + 2 cos(x) ? or\[\int\limits_{}^{}\frac{ dx }{ 2(1+\cos x) }?\] One approach that might work and is worth investigating: Mult. numerator and denom. of this integrand by the conjugage of (1+cos x); that is, mult. by (1-cos x).

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

does it help to know that \[\cos(x)+\sqrt{3}\sin(x)= 2\sin(\frac{\pi}{6}-x)\] or even \[-2\sin(x-\frac{\pi}{6})\]

OpenStudy (misty1212):

oops i think i am off by a signe

OpenStudy (misty1212):

no, just made a typo \[cos(x)-\sqrt3\sin(x)=-2\sin(x-\frac{\pi}{6})\]

OpenStudy (eliesaab):

Building on the above one can write that \[ \frac{1}{-2 \sqrt{3} \sin (x)+2 \cos (x)+2}=\frac{1}{2 \left(1-2 \sin \left(x-\frac{\pi }{6}\right)\right)} \]

OpenStudy (eliesaab):

to check you answer, \[ \int \frac{1}{2 \left(1-2 \sin \left(x-\frac{\pi }{6}\right)\right)} \, dx=\frac{\tanh ^{-1}\left(\frac{\tan \left(\frac{1}{12} (\pi -6 x)\right)+2}{\sqrt{3}}\right)}{\sqrt{3}} \]

OpenStudy (eliesaab):

The above gives a step by step solution